# Optimally-Weighted Maximum Mean Discrepancy Framework for Continual Learning

### v1

@@ 4.4 Ablation Study
In this section, we perform an ablation study to evaluate the individual contributions of the three terms—$\alpha$-term, $\beta$-term, and $\gamma$-term—in our proposed method. The aim is to investigate how each component influences the overall performance and identify the most critical factors for improving continual learning.
We systematically test different combinations of these terms in the loss function. These experiments are conducted using a memory buffer size of 500 on the CIFAR-10 dataset. The results are summarized in the Table LABEL:tab:Ablation_Study below:
Table 7: Ablation study results on CIFAR-10 with buffer size 500
Term Class-IL Task-IL
base 19.62±0.33 61.02±0.24
$\alpha$ 70.26±0.33 93.31±0.18
$\beta$ 66.04±2.54 94.18±0.29
$\gamma$ 19.62±0.04 84.56±4.84
$\alpha+\beta$ 73.88±1.16 94.44±0.41
$\alpha+\gamma$ 73.53±1.21 93.81±0.31
$\beta+\gamma$ 69.53±1.33 94.99±0.09
$\alpha+\beta+\gamma$ 75.29±0.75 94.94±0.41
Base Model. This variant only includes the standard loss function without any of the additional terms ($\alpha$-term, $\beta$-term, or $\gamma$-term). This helps establish a baseline performance for comparison. The base model performs very poorly because it does not incorporate any continual learning strategies, resulting in catastrophic forgetting.
$\alpha$-term, $\beta$-term, and $\gamma$-term. In this configuration, we evaluate the individual contributions of the $\alpha$-term (logits matching), $\beta$-term (label replay), and $\gamma$-term (teacher-student transfer). The results show that the inclusion of only the $\alpha$-term or the $\beta$-term, both of which involve using the memory buffer, already yields good performance. However, when only the $\gamma$-term is used, the performance is poor. This is because the teacher-student method relies on a strong teacher network, and in the absence of the memory buffer, the teacher also suffers from catastrophic forgetting.
Combination of two terms. When two terms are combined, the results are better than when any single term is used individually. Specifically, combining the $\alpha$-term and $\beta$-term (logits matching and label replay) leads to improved performance, as both terms utilize the memory buffer to preserve past knowledge. Similarly, combining the $\alpha$-term and $\gamma$-term (logits matching and teacher-student transfer) also results in better performance compared to using either term alone, as it ensures both alignment with previous experiences and the knowledge transfer from the teacher network. This demonstrates the complementary nature of these terms in mitigating catastrophic forgetting.
Full Model. This final configuration includes all three terms ($\alpha$, $\beta$, and $\gamma$) as in our proposed method. It represents the complete approach and serves as the optimal model for comparison.

### v4

@@ 4.5 Ablation Study
To analyze the effectiveness of multi-layer feature matching and adaptive regularization, we conduct an ablation study by selectively applying MMD to different layers of the backbone. Specifically, we compare models that use only a subset of layers (e.g., the first few or last few) against the full model, and we evaluate the impact of the Adaptive Regularization Optimization Term (AROT). This experiment is performed on CIFAR-10 with a buffer size of 500.
Table 8 shows that when AROT is enabled, using three layers for MMD regularization yields better performance than using only two, demonstrating that a broader feature alignment contributes to more effective continual learning. Notably, aligning the first three layers achieves superior results compared to the last three layers, indicating that shallow feature matching plays a more significant role in mitigating forgetting. This aligns with the observed adaptive weight trends, where lower-level layers receive higher importance as training progresses. Shallow layers capture fundamental structures that are transferable across tasks, whereas deeper layers extract task-specific features that may not generalize as effectively. Nevertheless, the best results are obtained when all layers are utilized with AROT, as this ensures a well-balanced integration of generalizable and task-specific representations.
When AROT is not used, performance degrades across all layer configurations, underscoring the importance of adaptive weighting. Without AROT, models exhibit greater fluctuations in accuracy depending on the selected layers, as fixed weighting fails to optimally balance the contributions of different feature levels. The consistent improvement observed with AROT highlights its role in dynamically adjusting the influence of each layer, enabling better feature retention and mitigating catastrophic forgetting.
Table 8: Ablation study on CIFAR-10 with buffer size 500. "First $n$" indicates that only the first $n$ layers of the backbone are utilized for matching and MMD calculation, while "Last $n$" refers to using only the last $n$ layers. "ALL" indicates that all backbone layers are used. AROT denotes Adaptive Regularization Optimization Term.
Layers With AROT No AROT
Class-IL Task-IL Class-IL Task-IL
First 1 - - 73.94±0.96 94.38±0.38
First 2 74.57±0.57 94.41±0.61 74.04±0.59 94.56±0.44
First 3 74.77±0.99 94.39±0.54 73.25±0.72 94.42±0.17
Last 3 74.80±0.38 94.44±0.06 74.46±0.79 94.42±0.45
Last 2 74.36±0.87 94.50±0.29 74.06±0.79 94.38±0.06
Last 1 - - 74.87±1.04 94.30±0.38
ALL 75.29±0.75 94.94±0.41 74.69±1.07 94.51±0.24